Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 120,000 items were shipped to customers using 2,300 direct labor-hours.

The company incurred a total of $7,360 in variable overhead costs. According to the company’s standards, 0.02 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.25 per direct labor-hour.

Required: 1. What is the standard labor-hours allowed (SH) to ship 120,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 120,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1
Part-1
Standard Labour Hour Allowed = Quantity X Std Hour Required per Hour
120000X 0.02= 2400 Hour
Part-2 :
Standard Variable Overhead Cost Allowed = Standard Labour Hour X Standard Variable Overhead rate per Hour
=2400 HourX $3.25 =$7800
Part-3:
Variable Overhead Spending Variance= (Actual Hour X Std Variable OH Rate per Hour)- Actual Overhead cost
(2300 HourX $3.25)-7360=$115 Favourable
Part-4
Variable OH Rate Variance= ( Std Rate - Actual Rate) Actual Hour
=($3.25-$3.20)X 2300=$115 Favourable
Actual Rate= 7360/2300=$3.20
Variable OH Efficiency Variance= ( SHOur- Actual Hour) X Std Rate
( 2400 Hour- 2300 Hour) X $3.25=$325 Favourable
Let me know if any doubt Kindly mark with positive rating it would help me lot. Thanks in advance.